INFRARED MATRIX ISOLATION STUDY OF THE REACTION OF TiCl_{4} WITH LEWIS ACIDS; SPECTROSCOPIC CHARACTERIZATION OF Cl_{3}TiOCH_{3}

\underline{BRUCE S. AULT} and JENNIFER B. EVERHART, Department of Chemistry, University of Cincinnati, P. O. Box 210172, Cincinnati, OH 45221-0172.

The reaction of TiCl_{4} with NH_{3}, CH_{3}OH and other Lewis bases in twin jet and merged jet deposition systems followed by trapping in argon matrices at 14 K has been examined. With twin jet deposition, formation of the initial intermediate, a 1:1 complex, was observed in each case. In a merged jet reactor, the reaction of TiCl_{4} with CH_{3}OH proceeded very rapidly to a second intermediate, monomeric Cl_{3}TiOCH_{3}, which has been observed for the first time. Identification and characterization was carried out by extensive isotopic labeling, and comparison to density functional calculations. Increasing the temperature of the merged jet region to above 200deg C led to the complete destruction of this species, and formation of additional products. Recent results on these and related systems will be reported.